OpenStudy (anonymous):

Which of the following relations is a function? A. {(5, 2), (3, 2), (4, 2), (5, 1)} B. {(1, 2), (1, 3), (2, 1), (3, 4)} C. {(1, 3), (2, 6), (3, 8), (1, 7)} D. {(1, 4), (8, 9), (10, 11), (12, 12)}

OpenStudy (e.mccormick):

Do you know the definition of a function when it comes to x,y pairs?

OpenStudy (anonymous):

is a set of ordered pairs where each y is paired with exactly one x?

OpenStudy (anonymous):

wait no a x is paired with one y right

OpenStudy (e.mccormick):

The second. For any given x there is one and only one y. So if you have an x of say 42, and it maps to 12 in one pair and 14 in another pair, it is NOT a function. On the other hand, if you have an x of 42 and an x of 24 and they both map to a y of 12, that is just fine!

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

Would the answer be D?

OpenStudy (e.mccormick):

D looks good. =)

OpenStudy (anonymous):

Wow thanks!

OpenStudy (technodynamic):

Th reason this is, it function/relation can never have more than one of the same in its sets.

OpenStudy (technodynamic):

so it's D.

OpenStudy (e.mccormick):

Once you get the basic concept, these are not too hard. Just remember it is unique x values. Now, when you get to one-to-one things, it needs to be both unique x and unique y.

OpenStudy (anonymous):

OK thanks i have a other question i have no idea how to do.

OpenStudy (anonymous):

Is the equation y = x3 - 6 even, odd, or neither? A. even B. odd C. neither

OpenStudy (anonymous):

This i don't know how to do.

OpenStudy (technodynamic):

is that a cubed x? or 3x?

OpenStudy (anonymous):

yes cubed

OpenStudy (technodynamic):

if you're doing quadratic eqn then it would be -x3+y+6=0 or, if you're solving just that equation get -y on both sides.

OpenStudy (e.mccormick):

No, he is asking what type of function it is, even, odd, or neither. \(y = x^3 - 6\) Do you know what the even and odd tests are?

OpenStudy (anonymous):

Not really i don't know

OpenStudy (e.mccormick):

\(f(-x)\) and then see what would happen if you follow the rules of algebra. See, if you have \(f(x)=x^2\) then \(f(-x)= (-x)^2\) Well, the square gets rid of the - sign. Therefore \(f(x)=f(-x)\) and the function is even.

OpenStudy (e.mccormick):

If \(-f(x)=f(-x)\) it is odd. If \(f(x)=f(-x)\) it is even. All other cases it is neither.

OpenStudy (e.mccormick):

Some examples: http://www.purplemath.com/modules/fcnnot3.htm

OpenStudy (anonymous):

This is kinda hard but i think i got it.

OpenStudy (e.mccormick):

If the x to -x is confusing, use a different letter. See if you put a and -a in and if they match or are opposites. Oh, and \(-f(x)\) means multiply through with -1, so any function with a constant, like +2 on the end, can NEVER be odd.

OpenStudy (anonymous):

Would the equation be something like -f(x)=x3-6?

OpenStudy (e.mccormick):

Well, for the odd test, it is this: Test -f(x) \(f(x)=x^2+2\) \((-1)f(x)=(-1)(x^2+2)\) \(-f(x)=-x^2-2\) and test f(-x) \(f(-x)=(-x)^2+2\) \(f(-x)=x^2+2\) and see if they match \(-f(x)=f(-x)\implies x^2+2=-x^2-2\) Because the last is not tue, it is not odd.

OpenStudy (e.mccormick):

Usually people test even first because f(-x) is all it needs, then you reuse the result in odd... but that is when you have to show all your work.

OpenStudy (anonymous):

did you like simplify the x3 and 6? and then was x2+2?

OpenStudy (e.mccormick):

No. I just gave an example. Not doing your work for you. =)

OpenStudy (anonymous):

ohh ok XD

OpenStudy (e.mccormick):

np. With \(x^n\) the \(-x\) part is pretty easy. If n is an odd number, the - stays. If the n is an even number, the - goes by-by.

OpenStudy (anonymous):

=(-1)(x3-6) =-x3+6

OpenStudy (anonymous):

so it be odd?

OpenStudy (anonymous):

@InYourHead

OpenStudy (e.mccormick):

Like I said, if it has a constant at the end, it can not be odd because the sign of the constant changes.

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

Can you show me a example of one being neither ?

OpenStudy (anonymous):

i get odd now. :)

OpenStudy (e.mccormick):

Well, any with a variable to an odd power is not even, right? And any with a constant term is not odd, right? So all the ones with odd powers and a constant term are neither even nor odd.

OpenStudy (anonymous):

soo would this be neither ? :o

OpenStudy (e.mccormick):

Yah. Here is a graph of three examples of each type. https://www.desmos.com/calculator/taotr3tsip You can turn on and off any graph at any time by clicking the colored dot next to the equation. Notice how the evens are all mirroed right to left, the odds are mirroed from Q1 to Q3 and Q2 to Q4 and the nones are not mirrors at all.

OpenStudy (anonymous):

one questino in order to be negative i need a even power on a variable

OpenStudy (anonymous):

thats a cool web site

OpenStudy (e.mccormick):

negative? You mean odd?

OpenStudy (anonymous):

yes odd*

OpenStudy (e.mccormick):

Even must have only even powers, may or may not have constants. Odd must only have odd powers, must not have costants.

OpenStudy (anonymous):

ok i get everything now thanks soo much

OpenStudy (e.mccormick):

Yah, looking at a graph can really help at times. It gives you something to focus on with visual memory, which is the biggest part of sensory memory. =)